OpenStudy (anonymous):

Sodium reacts with chlorine gas to form sodium chloride. Write a balanced chemical equation for the reaction, and find the mass of chlorine gas that will react with 96.6 g of sodium.

OpenStudy (anonymous):

2Na + Cl2 --> 2 NaCl

OpenStudy (anonymous):

\[96.6 g Na (\frac{1 mole Na}{22.99g Na})(\frac{1 mole Cl_2}{2 mole Na})(\frac{70.9 g Cl_2}{1 mole Cl_2})=149.0g Cl_2\]